To live comfortably in retirement, you decide you will need to save $2 million by the time you are 65 (you are 30 years old today). You will start a new retirement savings account today and contribute the same amount of money on every birthday up to and including your 65th birthday. Using TVM principles, how much must you set aside each year to make sure that you hit your target goal if the interest rate is 5%? What flaws might exist in your calculations, and what variables could lead to different outcomes? What actions could you take ensure you reach your target goal?
Answer #1

Amount to be save = 2,000,000
N = 35
r =5%
Future Value Amount = PMT * ( (1+r)n -1)/r + (PMT at age of 65)
Amount = PMT *((1.05)35-1)/0.05 + PMT
2,000,000 = 91.32 * PMT
Each Year PMT = 21,900.93

The flaw in the calculation is it does not include the effect of inflation and also interest rate doesn't remain constant it keeps on fluctuating in real life scenario .So including inflation would yield different results and the amount to be saved per year will increase.

Real interest rate should be calculated using inflation rates as base and the discounting factor will decrease , so payments per year has to be increased to achieve real life goals.